OpenStudy (anonymous):

How do you integrate this using substitution ? ∫sin (3pi/2 + pi/4)dx

OpenStudy (lgbasallote):

\[\large\int \sin (\frac{3\pi}{2} +\frac{\pi}{4})dx?\]

OpenStudy (mimi_x3):

I think it's missing a \(x\)

OpenStudy (anonymous):

yea thats the question

OpenStudy (anonymous):

Would the answer be sqrt(2)/2 x?

OpenStudy (anonymous):

the answer at the back is -2/3pi cos (3pi/2 + pi/4) +c

OpenStudy (mimi_x3):

Are you sure that it's \[\int sin\left(\frac{3\pi}{2}+\frac{\pi}{4}\right) dx\]?? I think that it's missing a \(x\) somewhere

OpenStudy (anonymous):

yea thats the question

OpenStudy (anonymous):

Ok so the answer is -sqrt(2)/2 x

OpenStudy (anonymous):

Just simplify sin ( (3pi/2) + (pi/4))

OpenStudy (anonymous):

sin ((3pi/2) + (pi/4)) = -sin(pi/4)

OpenStudy (anonymous):

Then if you remember your exact values, sin(pi/4) = sqrt(2)/2

OpenStudy (anonymous):

So -sin(pi/4) = -sqrt(2)/2

OpenStudy (anonymous):

Then integrate sqrt(2)/2

OpenStudy (anonymous):

I mean integrate -sqrt(2)/2 with respect to x

OpenStudy (anonymous):

You will get integral(sin ( (3pi/2) + (pi/4))) = -sqrt(2)/2 x

OpenStudy (anonymous):

thank you :)